# Enumerative Geometry of Quantum Periods

### v1

@@ 5. Open-log correspondence for smooth log Calabi-Yau pairs
In this section, we prove a higher genus open-log correspondence for smooth log Calabi-Yau pairs $(\mathbb{P},E)$. We first show a correspondence between higher genus two-pointed log invariants and local invariants in a similar approach as in [GRZ]. Then assuming $\mathbb{P}$ is toric, we use the Topological Vertex [AKMV] to relate the log invariants to open invariants of $K_{\mathbb{P}}$ in all genera.
@@
In order to state our results, we first establish some notation. Let $w\in\mathbb{Z}_{\geq 0}$, and $\vec{k}=(k_{1},\ldots,k_{n})\in\mathbb{Z}_{\geq 0}^{n}$ be a vector of non-negative integers such that $\sum_{j=1}^{n}jk_{j}=w$. Let $|\vec{k}|$ be the number of non-zero entries of $\vec{k}$. We denote $O_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ to be the genus-$g$, winding-$w$, open Gromov-Witten invariant of an outer Aganagic-Vafa brane $L\subset K_{\mathbb{P}}$ in framing-0, in curve class $\beta+w\beta_{0}$ with $\beta_{0}\in H_{2}(K_{\mathbb{P}},L)$, and in winding profile $\vec{k}$ if it has $k_{j}$ boundary components of winding $j$. The invariants $O_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ can be defined using stable relative maps, and we refer to [LLLZ] [FL] for a more detailed definition. We write $n_{g}^{open}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ to be the corresponding open-BPS invariant defined from open multiple cover formulas for $O_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ [MV]. We will frequently consider winding-$w$ invariants with 1 boundary component in which case the winding multiplicity vector is $\vec{k}=(0,\ldots,1,0,\ldots)$ has a single 1 in the $w$-th position corresponding to a single hole with winding $w$. As Gromov-Witten invariants and BPS invariants are uniquely determined by each other via the multiple-cover formulas, the $n^{open}_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ can be defined by the $O_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$.
@@
Let $\pi:\hat{\mathbb{P}}\rightarrow\mathbb{P}$ be the blow up of $\mathbb{P}$ at a point with exceptional curve $C$. In Section 5.1, we first review the genus-0 open-log correspondence for smooth log Calabi-Yau pairs $(\mathbb{P},E)$ established in [GRZ]. In Section 5.2, we relate higher genus two-pointed log invariants $R_{g,(1,p)}(\mathbb{P}(\log E),\beta)$ with Gopakumar-Vafa invariants of local $\widehat{\mathbb{P}}$ in Theorem 5.2. In Section 5.3, additionally assuming $\mathbb{P}$ is toric, we use the Topological Vertex [AKMV] to relate $O_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ to closed invariants of $K_{\widehat{\mathbb{P}}}$ in curve class $\pi^{*}\beta-wC$ in Theorem 5.3, and then specialize it to winding-1 in Corollary 5.4 to obtain an open-closed equality of BPS invariants that generalizes Theorem 1.1, [LLW] to all genera. In Section 5.5, we prove in Theorem 5.8 an open-log correspondence in all genera that relates two-pointed log invariants $R_{g,(1,p)}(\mathbb{P}(\log E),\beta)$ to winding-1, 1-boundary component, open-BPS invariants $n_{g}^{open}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$ via a discrepancy quantified by stationary Gromov-Witten invariants of $E$, and $R_{h,(1,p)}(\mathbb{P}(\log E),\beta)$ for $h<g$. We provide a computational check of Theorem 5.8 in Example 5.10 for $\mathbb{P}=\mathbb{C}\mathbb{P}^{2}$. In Section 5.6, we provide a computational check of Corollary 5.4 using the topological vertex with the relevant open and closed BPS invariants listed in Table 1.
@@ 5.1. The $g=0$ open-log correspondence
When $\mathbb{P}=\mathbb{C}\mathbb{P}^{2}$, the above invariants are the coefficients of the open mirror map $M(Q)$ of an outer Aganagic-Vafa brane $L\subset K_{\mathbb{C}\mathbb{P}^{2}}$ in framing-0,
@@
Theorem 5.1 implies that the mirror dual Landau-Ginzburg superpotential is equivalent to the open mirror map $M(Q)$. Its proof uses results from [Gra2], [vGGR], [LLW], and Corollary 6.6 of [GRZ] to relate two-pointed log invariants $R_{0,(1,\beta\cdot E-1)}$ to open invariants of $K_{\mathbb{P}}$.
@@ 5.2. Higher genus log/Gopakumar-Vafa correspondence
By Corollary 6.6 of [GRZ], the above generating series becomes,
@@
where $R_{g}(\widehat{\mathbb{P}}(\log\pi^{*}E-C),\pi^{*}\beta-C)$ are genus-$g$, maximal tangency invariants of the log Calabi-Yau pair (^$(\widehat{\mathbb{P}},\pi^{*}E-C)$ with $\lambda_{g}$-insertion (defined in Section 4, [GRZ]), and $N(g,p)$ is the genus-$g$, local relative invariant of $\mathbb{P}^{1}(\log\infty)$ with tangency order $p$ computed in Theorem 5.1, [BP], and is given by the $\hbar^{2g}$-coefficient of $\frac{(-1)^{p+1}}{p}\frac{i\hbar}{q^{p/2}-q^{-p/2}}$.
@@
Applying the $g\geq 0$ log-local principle (combining Propositions 3.1 and 3.4 of [BFGW]) to $R_{g,(\beta\cdot E-1)}(\widehat{\mathbb{P}}(\log\pi^{*}E-C),\pi^{*}\beta-C)$ in (5.1), it becomes,
@@
5.3. Winding-$w$ open/closed correspondence
For this section, suppose that $\mathbb{P}$ is toric and $\pi:\widehat{\mathbb{P}}\rightarrow\mathbb{P}$ is a toric blow up at a point. Then $K_{\mathbb{P}},K_{\widehat{\mathbb{P}}}$ are toric Calabi-Yau 3-folds. We use the Topological Vertex [AKMV] to derive an equality relating the winding-$w$, open invariants $O_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ of an outer AV-brane $L\subset K_{\mathbb{P}}$ in winding profile $\vec{k}$ to the closed invariants $N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)$. The work of [LLLZ] [LLZ1] [LLZ2] [MOOP] show that calculations with the Topological Vertex compute the open Gromov-Witten invariants $O_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$. We refer to [Kon] [LLLZ] [AKMV] for more definitions and details on the topological vertex.
The toric diagram of a local toric Fano surface $\mathbb{P}$ is given by a convex $k$-gon with a non-compact edge emanating from each vertex. We refer to Figure 5.1 for the relevant toric diagrams. Let $\{v_{i}\}_{i=1}^{k}$ be the torus-fixed points, and $\{\ell_{i}\}_{i=1}^{k}$ be the compact 1-dimensional torus orbits of the toric diagram of local $\mathbb{P}$. Let $t_{i}$ be the Kähler parameter of $\ell_{i}$, and define the variable $Q^{\ell_{i}}:=e^{-t_{i}}$. For each $\ell_{i}$, we assign a variable $\lambda_{i}$ that runs over all partitions of arbitrary length. We refer to a partition $\lambda$ and its $U(N$)-representation interchangeably. By the topological vertex, the partition function of local $\mathbb{P}$ is given by,
@@ 5.3. Winding-$w$ open/closed correspondence
where the vertex functions $C_{\lambda_{1}\lambda_{2}\lambda_{3}}$ are defined in Definition 2.2, [Kon].
The toric diagram of local $\widehat{\mathbb{P}}$ has an additional compact 1-dimensional torus orbit in the polygon corresponding to the exceptional curve $C\cong\mathbb{P}^{1}$, with Kähler parameter $t_{k+1}$ and $Q^{C}:=e^{-t_{k+1}}$. We label the partition of $C$ to be $\lambda_{k+1}$. The partition function of local $\widehat{\mathbb{P}}$ is given by,
@@
In addition, we define $Z_{\widehat{\mathbb{P}}}|_{\pi^{*}\beta-wC}$ to be,
$Z_{\widehat{\mathbb{P}}}(Q^{\ell_{i}},Q^{C})|_{\pi^{*}\beta-wC}:=Z_{\widehat{\mathbb{P}}}(Q^{\ell_{i}},Q^{-C})|_{|\lambda_{k+1}|=w}$
As the normal bundle of $C$ in $K_{\widehat{\mathbb{P}}}$ is $\mathcal{O}_{\mathbb{P}^{1}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{1}}(-1)$, denote $W$ to be the space obtained by flopping along $C\subset K_{\widehat{\mathbb{P}}}$. We write $E$ for the exceptional curve of $W$, with Kähler parameter $t^{\prime}_{k+1}$ and $Q^{E}:=e^{-t^{\prime}_{k+1}}$. We label the partition of $E$ to be $\lambda^{\prime}_{k+1}$. The partition function of $W$ is given by,
@@
Let $V$ be an $U(N)$-matrix with $N$ suitably large, representing the monodromy of a flat connection on $L$. Let $Tr_{\nu}V$ be the trace of $V$ in a $U(N)$-representation given by a partition $\nu$. The open free energy of a single outer AV-brane $L\subset K_{\mathbb{P}}$ is given by,
@@
For a fixed partition or representation $\nu$ attached to an outer AV-brane, we define,
@@
We have the following theorem.
@@
Let $\mathbb{P}$ be a toric del Pezzo surface, and $\beta\in\textup{NE}(\mathbb{P})$ an effective curve class. Let $\pi:\widehat{\mathbb{P}}\rightarrow\mathbb{P}$ be the toric blow up at a point with exceptional curve $C$. Let $K_{\mathbb{P}},K_{\widehat{\mathbb{P}}}$ be the toric canonical bundles. We have the equality of winding-$w$ open and closed partition functions,
$\sum_{|\lambda|=w}Z_{\mathbb{P}}(V)|_{\lambda}Z_{\mathbb{C}^{3}}(V)|_{\lambda^{T}}=Z_{\widehat{\mathbb{P}}}|_{\pi^{*}\beta-wC}$
where $Z_{\mathbb{C}^{3}}(V)|_{\lambda^{T}}$ is the winding-$w$, open partition function of an outer AV-brane in $\mathbb{C}^{3}$ in partition $\lambda^{T}$.
Equivalently, writing $\beta=c_{1}\ell_{1}+\ldots c_{k}\ell_{k}$, the invariant $N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)$ is the $\hbar^{2g}$-coefficient of the expression,
@@
$\sum_{\begin{subarray}{c}l\geq 0,l|w,\\ (w_{1},\ldots,w_{l})\in\mathbb{Z}^{l}_{\geq 0},\\ 1\leq j\leq l,|(\lambda^{\prime}_{k+1})_{j}|=w_{j}\end{subarray}}\prod_{i=1}^{l}\sum_{\begin{subarray}{c}1\leq a\leq k,\\ |\lambda_{ia}|=c_{ia},\\ c_{ia}\in\mathbb{Z}_{\geq 0}\end{subarray}}\left[\frac{(-1)^{l+1+\sum_{a}|\lambda_{ia}|}}{l}q^{\sum_{a}\kappa_{\lambda_{ia}}}\left(\prod_{a=1}^{k-1}C_{\lambda^{T}_{ia}\lambda_{(i+1)a}\varnothing}\right)C_{\lambda^{T}_{ka}\lambda_{1a}(\lambda^{\prime}_{k+1})_{i}}C_{(\lambda^{{}^{\prime}T}_{k+1})_{i}\varnothing\varnothing}\right]$
where $q=e^{i\hbar}$, $(w_{1},\ldots,w_{l})\in\mathbb{Z}_{\geq 0}^{l}$ is a partition of $w$, $(\lambda^{\prime}_{k+1})_{j}$ is a partition of size $w_{j}$ for $1\leq j\leq l$, $(c_{ia})\in\mathbb{Z}_{\geq 0}^{l}$ is a partition of $c_{a}$ for $1\leq a\leq k$.
Proof of Theorem 5.3.
By the gluing formula for the topological vertex or Proposition 5.8, [LLLZ], which is shown by the degeneration formula for relative Gromov-Witten invariants [Li], we can express $Z_{W}$ in terms of winding-$w$, open Gromov-Witten invariants of an outer AV-brane $L\subset K_{\mathbb{P}}$ and the open invariants of an outer AV-brane in $\mathbb{C}^{3}$. We have,
$Z_{W}(Q^{\ell_{i}},Q^{E})$ $=\sum_{\lambda^{\prime}_{k+1}}Z_{\mathbb{P}}(Q^{\ell_{i}},V)|_{\lambda^{\prime}_{k+1}}Z_{\mathbb{C}^{3}}(V)|_{\lambda^{{}^{\prime}T}_{k+1}}$
where $Z_{\mathbb{P}}(Q^{\ell_{i}},V)|_{\lambda^{\prime}_{k+1}}$ is the partition function of winding-$w$, open Gromov-Witten invariants of an outer AV-brane in $K_{\mathbb{P}}$ in representation $\lambda^{\prime}_{k+1}$, and $Z_{\mathbb{C}^{3}}(V)|_{\lambda^{{}^{\prime}T}_{k+1}}$ is the partition function of winding-$w$, open Gromov-Witten invariants of an outer AV-brane in $\mathbb{C}^{3}$ in representation $\lambda^{{}^{\prime}T}_{k+1}$. In particular, from Equation (6.9), [BBvG], we have,
$Z_{\mathbb{C}^{3}}(V)|_{\lambda^{{}^{\prime}T}_{k+1}}:=C_{\lambda^{\prime}_{k+1}\varnothing\varnothing}=q^{\kappa_{\lambda^{\prime}_{k+1}}/2}\sum_{\eta}s_{\lambda^{{}^{\prime}T}_{k+1}/\eta}(q^{\rho})$

### v2

@@ 5. Open-log correspondence for smooth log Calabi-Yau pairs
In this section, we prove a higher genus open-log correspondence for smooth log Calabi-Yau pairs $(\mathbb{P},E)$. We first show a correspondence between higher genus two-pointed log invariants and local invariants in a similar approach as in [GRZ]. Then assuming $\mathbb{P}$ is toric, we use the Topological Vertex [AKMV] to relate the log invariants of $(\mathbb{P},E)$ to open invariants of $K_{\mathbb{P}}$ in all-genus.
@@
In order to state our results, we first establish some notation. Let $w\in\mathbb{Z}_{\geq 0}$, and $\vec{k}=(k_{1},\ldots,k_{n})\in\mathbb{Z}_{\geq 0}^{n}$ be a vector of non-negative integers such that $\sum_{j=1}^{n}jk_{j}=w$. Let $|\vec{k}|$ be the number of non-zero entries of $\vec{k}$. We denote $O_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ to be the genus-$g$, winding-$w$, open Gromov-Witten invariant of an outer Aganagic-Vafa brane $L\subset K_{\mathbb{P}}$ in framing-0, in curve class $\beta+w\beta_{0}$ with $\beta_{0}\in H_{2}(K_{\mathbb{P}},L)$, and in winding profile $\vec{k}$ if it has $k_{j}$ boundary components of winding $j$. The invariants $O_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ can be defined using stable relative maps, and we refer to [LLLZ] [FL] for a more detailed definition. We write $n_{g}^{open}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ to be the corresponding open-BPS invariant defined from open multiple cover formulas for $O_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ [MV]. As Gromov-Witten invariants and BPS invariants are uniquely determined by each other via the multiple-cover formulas, the $n^{open}_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ can be defined by the $O_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$.
@@
Let $\pi:\hat{\mathbb{P}}\rightarrow\mathbb{P}$ be the blow up of $\mathbb{P}$ at a point with exceptional curve $C$. In Section 5.1, we first review the genus-0 open-log correspondence for smooth log Calabi-Yau pairs $(\mathbb{P},E)$ established in [GRZ]. In Section 5.2, we relate higher genus two-pointed log invariants $R_{g,(1,p)}(\mathbb{P}(\log E),\beta)$ with Gopakumar-Vafa invariants of local $\widehat{\mathbb{P}}$ in Theorem 5.2. In Section 5.3, additionally assuming $\mathbb{P}$ is toric and for winding $w>0$, we use the Topological Vertex [AKMV] to relate Gopakumar-Vafa invariants of local $\widehat{\mathbb{P}}$ in curve class $\pi^{*}\beta-wC$ to certain LMOV invariants of $L\subset K_{\mathbb{P}}$ in representation $(w)$ that is a Young Tableau with a single row of $w$ boxes in Theorem 5.3. LMOV invariants are related to the open-BPS invariants $n_{g}^{open}$ by Equation 5.8. We then specialize Theorem 5.3 to $w=1$ in Corollary 5.5 that generalizes [LLW], Theorem 1.1 to all-genus, and to $w=2$ in Corollary 5.7. We conjecture a higher-winding open-closed relation involving LMOV invariants in Conjecture 5.8.
In Section 5.4, we prove an open-log correspondence in all-genus in Theorem 5.13 which relates two-pointed log invariants $R_{g,(1,p)}(\mathbb{P}(\log E),\beta)$ to winding-1, 1-boundary component, open-BPS invariants $n_{g}^{open}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$ via a discrepancy quantified by stationary Gromov-Witten invariants of $E$, and $R_{h,(1,p)}(\mathbb{P}(\log E),\beta)$ for $h<g$. We provide a computational check of Theorem 5.13 in Example 5.14 for $\mathbb{P}=\mathbb{P}^{2}$. In Section 5.5, we provide a computational check of Corollary 5.5 using the Topological Vertex with the relevant open and closed BPS invariants listed in Table 1.
@@ 5.1. The $g=0$ open-log correspondence
When $\mathbb{P}=\mathbb{P}^{2}$, the above invariants are the coefficients of the open mirror map $M(Q)$ of an outer Aganagic-Vafa brane $L\subset K_{\mathbb{P}^{2}}$ in framing-0,
@@
Theorem 5.1 implies that the mirror dual Landau-Ginzburg superpotential is equivalent to the open mirror map $M(Q)$. Its proof uses results from [Gra2] [vGGR] [LLW] and [GRZ], Corollary 6.6 to relate two-pointed log invariants $R_{0,(1,\beta\cdot E-1)}$ to open invariants of $K_{\mathbb{P}}$.
@@ 5.2. Higher genus log/Gopakumar-Vafa correspondence
By [GRZ], Corollary 6.6, the above generating series becomes,
@@
where $R_{g}(\widehat{\mathbb{P}}(\log\pi^{*}E-C),\pi^{*}\beta-C)$ are genus-$g$, maximal tangency invariants of the log Calabi-Yau pair (^$(\widehat{\mathbb{P}},\pi^{*}E-C)$ with $\lambda_{g}$-insertion (defined in [GRZ], Section 4), and $N(g,p)$ is the genus-$g$, local relative invariant of $\mathbb{P}^{1}(\log\infty)$ with tangency order $p$ computed in [BP], Theorem 5.1 and is given by the $\hbar^{2g}$-coefficient of $\frac{(-1)^{p+1}}{p}\frac{i\hbar}{q^{p/2}-q^{-p/2}}$.
@@
Applying the $g\geq 0$ log-local principle (combining [BFGW], Propositions 3.1 and 3.4) to $R_{g,(\beta\cdot E-1)}(\widehat{\mathbb{P}}(\log\pi^{*}E-C),\pi^{*}\beta-C)$ in (5.1), it becomes,
@@
5.3. Open/closed correspondence
For this section, suppose that $\mathbb{P}$ is toric and $\pi:\widehat{\mathbb{P}}\rightarrow\mathbb{P}$ is a toric blow up at a point. Then $K_{\mathbb{P}},K_{\widehat{\mathbb{P}}}$ are toric Calabi-Yau 3-folds. For $w>0$, we use the Topological Vertex [AKMV] to derive an equality relating the genus-$g$, degree-$d$ LMOV invariant $N^{\mathrm{LMOV}}_{g,(w)}(K_{\mathbb{P}}/L,\beta)$ of an outer AV-brane $L\subset K_{\mathbb{P}}$ with Young Tableau given by a single row $(w)$ of length $w$ boxes to the closed invariants $N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)$. By Equation 5.8 and open multiple cover formulas [MV], this relates $O_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$, where $\vec{k}$ satisfies $\sum_{j}jk_{j}=w$, to $N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)$. The work of [LLLZ] [LLZ1] [LLZ2] [MOOP] show that calculations with the Topological Vertex [AKMV] compute the open Gromov-Witten invariants $O_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ as defined with stable relative maps. We refer to [Kon] [LLLZ] [AKMV] for more definitions and details on the Topological Vertex.
The toric diagram of a local toric Fano surface $\mathbb{P}$ is given by a convex $k$-gon with a non-compact edge emanating from each vertex. We refer to Figure 5.1 for the relevant toric diagrams. Let $\{v_{i}\}_{i=1}^{k}$ be the torus-fixed points, and $\{\ell_{i}\}_{i=1}^{k}$ be the compact 1-dimensional torus orbits of the toric diagram of local $\mathbb{P}$. Let $t_{i}$ be the Kähler parameter of $\ell_{i}$, and define the variable $Q^{\ell_{i}}:=e^{-t_{i}}$. For each $\ell_{i}$, we assign a variable $\lambda_{i}$ that runs over all partitions or Young Tableaus of arbitrary length. We refer to a partition $\lambda$ and its $U(N$)-representation interchangeably. By the Topological Vertex, the partition function of local $\mathbb{P}$ is given by,
@@ 5.3. Open/closed correspondence
where the vertex functions $C_{\lambda_{1}\lambda_{2}\lambda_{3}}$ are defined in [Kon], Definition 2.2. We will write a partition $\lambda$ as a finite sequence of non-increasing integers $\lambda_{1}\geq\lambda_{2}\geq\ldots$.
The toric diagram of local $\widehat{\mathbb{P}}$ has an additional compact 1-dimensional torus orbit in the polygon corresponding to the exceptional curve $C\cong\mathbb{P}^{1}$, with Kähler parameter $t_{k+1}$ and $Q^{C}:=e^{-t_{k+1}}$. We label the partition of $C$ as $\lambda_{k+1}$. The partition function of local $\widehat{\mathbb{P}}$ is given by,
@@
We define $Z_{\widehat{\mathbb{P}}}(Q^{\ell_{i}},Q^{-C})$ by replacing $Q^{C}$ with $Q^{-C}$ in $Z_{\widehat{\mathbb{P}}}(Q^{\ell_{i}},Q^{C})$.
As the normal bundle of $C$ in $K_{\widehat{\mathbb{P}}}$ is $\mathcal{O}_{\mathbb{P}^{1}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{1}}(-1)$, we denote $W$ to be the space obtained by flopping along $C\subset K_{\widehat{\mathbb{P}}}$. We write $E$ to be the corresponding exceptional curve of $W$, with Kähler parameter $t^{\prime}_{k+1}$ and $Q^{E}:=e^{-t^{\prime}_{k+1}}$. We label the partition of $E$ to be $\lambda^{\prime}_{k+1}$. The partition function of $W$ is given by,
@@
Let $V$ be an $U(N)$-matrix with $N$ suitably large, representing the monodromy of a flat connection on $L$. Let $Tr_{\nu}V$ be the trace of $V$ in a $U(N)$-representation given by a partition $\nu$. The open free energy of an outer AV-brane $L\subset K_{\mathbb{P}}$ is given by,
@@
For a fixed partition $\nu$ attached to an outer AV-brane, define,
@@
We write $Z_{\mathbb{P}}(Q^{\ell_{i}},V):=\sum_{\nu}Z_{\mathbb{P}}|_{\nu}Tr_{\nu}V$. Define the normalized open partition function $Z^{open}_{\mathbb{P}}(V):=\frac{Z_{\mathbb{P}}(Q^{\ell_{i}},V)}{Z_{\mathbb{P}}(Q^{\ell_{i}})}$, which is given by,
$Z^{open}_{\mathbb{P}}(V)=\mathrm{Exp}\left(\sum_{\nu}f_{\nu\varnothing\varnothing}(q,Q)Tr_{\nu}V\right)$
where $\mathrm{Exp}$ denotes the plethystic exponential.  [^7] We mention a minor typo in [AKMV], Equation 7.23. The subscript of $\widehat{f}$ should be $R^{\prime}_{i}$ not $R_{i}$. The $f_{\nu\varnothing\varnothing}(q,Q)$ are generating series related to LMOV invariants in representation $\nu$ of an outer AV-brane, and hence the $Z_{\nu}$ are expressible by LMOV invariants. Writing $Z^{open}_{\mathbb{P}}(V)=\sum_{\nu}Z_{\nu}Tr_{\nu}V$, we have $Z_{\mathbb{P}}|_{\nu}=Z_{\nu}Z_{\mathbb{P}}$.
Denote $N^{\mathrm{LMOV}}_{g,\lambda}(K_{\mathbb{P}}/L,\beta)$ to be the genus-$g$, LMOV invariant of an outer AV-brane $L\subset K_{\mathbb{P}}$ in degree $\beta$ and representation $\nu$. We write its generating series as,
(5.7) $\widehat{f}_{\nu\varnothing\varnothing}(q,Q)=\sum_{\begin{subarray}{c}g\geq 0,\\ \beta\in NE(\mathbb{P})\end{subarray}}N^{\mathrm{LMOV}}_{g,\nu}(K_{\mathbb{P}}/L,\beta)(q^{\frac{1}{2}}-q^{\frac{-1}{2}})^{2g}Q^{\beta}$
The $\widehat{f}_{\nu\varnothing\varnothing}$ are related to $f_{\nu\varnothing\varnothing}$ by [AKMV], Equation 7.23. The LMOV invariants are related to the open-BPS invariants by the linear transformation,
(5.8) $n_{g}^{open}(K_{\mathbb{P}}/L,\beta+\beta_{0},\vec{k})=\sum_{\nu}\chi_{\nu}(C(\vec{k}))N^{\mathrm{LMOV}}_{g,\nu}(K_{\mathbb{P}}/L,\beta)$
where $\chi_{\nu}(C(\vec{k}))$ is the character $\chi_{\nu}$ of the $U(N)$-representation $\nu$ evaluated on the conjugacy class $C(\vec{k})$ of the symmetric group $S_{w}$ determined by $\vec{k}$. The matrix $(\chi_{\nu}(C(\vec{k})))_{\nu,\vec{k}}$ is invertible over $\mathbb{Q}$ but not $\mathbb{Z}$. This suggests the LMOV invariants $N^{\mathrm{LMOV}}_{g,\nu}$ are in some sense more fundamental than the open-BPS invariants $n_{g}^{open}$. For more details on $Z^{open}_{\mathbb{P}}(V)$, LMOV invariants and their relationship with open-BPS invariants, we refer to [AKMV], Section 7.3.
Recall that $C\subset K_{\widehat{\mathbb{P}}}$ is the exceptional curve, $W$ is the flop of $K_{\widehat{\mathbb{P}}}$ along $C$, and $E$ is the resulting exceptional curve of $W$.
@@
Let $\pi:\widehat{\mathbb{P}}\rightarrow\mathbb{P}$ be the toric blow up at a point. Under the identification $Q^{-C}=Q^{E}$, we have the equality of generating series relating closed invariants $N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)$ to $\mathrm{LMOV}$ invariants $Z_{\nu}$ of an outer AV-brane $L\subset K_{\mathbb{P}}$,
@@
$\sum_{\begin{subarray}{c}g\geq 0,\\ \beta\in NE(\mathbb{P}),\\ w\in\mathbb{Z}_{>0}\end{subarray}}N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)\hbar^{2g-2}Q^{\pi^{*}\beta-wC}=\sum_{k\geq 1}\frac{(-1)^{k+1}}{k}\left(\sum_{\nu\neq\emptyset}(-1)^{|\nu|}Z_{\nu}C_{\nu^{T}\varnothing\varnothing}Q^{|\nu|E}\right)^{k}$
Hence, for $w\in\mathbb{Z}_{>0}$ the coefficient of $Q^{-wC}=Q^{wE}$ is given by,
$\sum_{j=1}^{w}\frac{(-1)^{j+1+w}}{j}\sum_{\begin{subarray}{c}\ell_{a}\in\mathbb{Z}_{\geq 1},\\ (\ell_{1},\ldots,\ell_{j}),\\ \sum_{a=1}^{j}\ell_{a}=w\end{subarray}}\prod_{\begin{subarray}{c}|\nu_{a}|=\ell_{a},\\ 1\leq a\leq j\end{subarray}}Z_{\nu_{a}}C_{\nu_{a}^{T}\varnothing\varnothing}$
Proof.